4. (2 pts) How much ATP is formed from aerobic metabolism of a 24:0 fatty acid?

5. (2 pts) Draw the following Fatty Acid: 24:2Δ5,9
Answer #1

Answer-

According to the given question-

We know that the beta-oxidation of fatty acids occurs in three stages inside the mitochondria–

The first stage – during this stage the molecule of Acetyl Co-A is removed from the fatty acids chain, here we have 24 carbon fatty acid chain of lignoceric acid, so we need 11 cycle, which leads to the production of 12 molecules of Acetyl Co A. during each cycle during the Acetyl Co A formation, two pairs of electrons as well as four Hydrogen are removed from the chain of fatty acid by dehydrogenases enzyme.

The second stage - during this stage, the Acetyl Co-A is oxidized into CO2 in TCA or citric acid cycle, inside the mitochondrial matrix, and also produces NADH and FADH2, a reduced electron carrier.

The third stage – during this stage, the NADH as well as FADH2 donate their electrons in the electron transport chain to oxygen, which reduces into the water and ATP formation takes place with the help of enzyme ATP synthetase.

So the equation for one cycle is –

lignoceric acid + CoA + FAD + NAD + + H2O ---- > Behenoyl Co A + Acetyl Co A + NADH + H+ + FADH2

So one molecule of lignoceric acid produces 12 molecules of acetyl-CoA

Overall equation will be-

lignoceric acid + 11 CoA + 11 FAD + 11 NAD + + 11 H2O ---- > 12 Acetyl-CoA + 11 FADH2 + 11 NADH + 11 H+

ATP formed from one molecule of –

1 NADH = 2.5 ATP and 1 FADH2 = 1.5 ATP , so Total = 4 ATP

So 4 ATP molecules are produced during each cycle to produce Acetyl CoA from the fatty acid chain.

Molecules of Acetyl CoA produced = 12

Molecules of ATP produced in this part = 11 * 4 = 44

So the overall equation for oxidation of lignoceric acid into 12 molecules of acetyl-CoA, following oxidative phosphorylation as well as electron transfers =

Lignoceriyl- Co A + 11 CoA + 11 O2 + 44 Pi + 44 ADP ---- > 12 Acetyl-CoA + 44 ATP + 11 H2O ----(equation 1)

During the TCA or citric acid cycle , the molecule of acetyl-CoA is oxidized into CO2 and H2O –

12 Acetyl-CoA + 24 O2 + 120 Pi + 120ADP ----- > 12 CoA + 120 ATP + 24 CO2 + 24 H2O. -- - (equation 2)

For 24 carbon saturated fatty acid- Number of pass = 11 (1 Acetyl Co A in each pass) Number of Acetyl Co-A formed = 12.

During each pass 1 molecule of FADH2 and 1 molecule of NADH is produced (So 1 FADH2 = 1.5 ATP, while 1 NADH = 2.5 ATP, thus in each pass = 4 ATP)

Total pass = 11 , Number of NADH = 11 and FADH2 = 11, Total ATP = 11 * 4 = 44 ATP.

Acetyl-CoA produced due to oxidation of fatty acids is oxidized into H2O and CO2 in citric acid cycle. Each acetyl Co A during citric acid cycle produces 10 ATP (3 NADH = 2.5 * 3 = 7.5 ATP, 1 FADH2 = 1.5 ATP and 1 GTP = 1 ATP , total = 7.5 + 1.5 + 1 = 10 ATP) ,

Here we have 12 molecules of Acetyl Co A , so total ATP = 12 * 10 = 120 ATP .

When we combine the number of ATP produced during the first part i.e. 44 ATP and during the TCA cycle (120 ATP)

Then total ATP = 120 + 44 = 164 ATP,

But 2 molecules of ATP are consumed during the activation of lignoceric acid into Lignoceriyl- Co A, thus Net ATP = 164 - 2 = 162 ATP.

Net ATP = 162 ATP.

Dear student, according to HOMEWORKLIB RULES I am only supposed to answer the first question. please post the question separately, if you want the answer. Thanks.